OpenStudy (anonymous):

(f-x)(g-x) = x+3-x-3) =?

OpenStudy (angela210793):

0

OpenStudy (anonymous):

Lol, simple example of 2+1-2-1?

OpenStudy (angela210793):

x-x=0 3-3=0 100-100=0 300000-300000=0